# MM-Sonate: Multimodal Controllable Audio-Video Generation with Zero-Shot Voice Cloning]Kling Team, Kuaishou Technology

### v1

@@ 1 Introduction
The ultimate goal of generative world simulation is to replicate the multisensory nature of physical reality, necessitating the synchronous synthesis of visual and acoustic signals. While recent text-to-video diffusion models have achieved remarkable visual fidelity, the integration of high-quality, synchronized audio remains a significant challenge. Early approaches typically treated audio generation as a post-processing step, employing cascaded systems where audio models are conditioned on pre-generated video frames [1, 2, 3, 4, 5, 6]. However, this decoupled paradigm fundamentally struggles with fine-grained temporal alignment, frequently resulting in significant asynchrony between video and audio streams.
To address the alignment problem, research has shifted toward unified architectures that generate audio and video within a single model. Pioneering works such as MM-Diffusion [7] and JavisDiT [8] explored joint modeling, while state-of-the-art models like Ovi [9] achieve impressive synchronization through symmetric dual-backbone designs. Despite these advances, a critical limitation persists: existing unified models lack fine-grained controllability over the acoustic modality, particularly in speech generation. For instance, while Ovi can generate semantically relevant sounds, it cannot control the specific identity (timbre) of a speaker or clone a voice from a reference clip. Consequently, current unified video generators cannot yet function as high-fidelity, personalized speech synthesis, limiting their application in scenarios requiring consistent character identity or dubbing.
@@
11footnotetext: Equal contribution.22footnotetext: “Sonate” is derived from the musical term “Sonata”, symbolizing the model’s comprehensive capabilities in audio generation.
@@ 2.1 Audio-Video Joint Generation
Audio-Video Joint Generation aims to simultaneously synthesize visual frames and acoustic signals, typically conditioned on text descriptions, ensuring both semantic consistency and temporal synchronization across modalities within a unified generation process. Pioneering efforts like MM-Diffusion [7] introduced this task using coupled U-Net backbones but were often constrained by small-scale datasets (e.g., roughly 10 hours), which limited their generalization capabilities. Following the architectural shift in unimodal generation, recent works have predominantly adopted Diffusion Transformers (DiTs). These approaches generally fall into two architectural paradigms: unified or coupled streams. For instance, UniForm [10] employs a unified single-tower architecture to process concatenated audio-video tokens, whereas AV-DiT [11] adapts a pre-trained image DiT with lightweight adapters to handle multimodal signals. A central challenge in this field remains precise spatio-temporal synchronization. To address this, JavisDiT [8] introduces a hierarchical prior, while Ovi [9] enforces tight coupling through a symmetric twin-backbone design. Similarly, SyncFlow [12] utilizes a dual-DiT architecture with Rectified Flow Matching to enhance alignment stability. Apart from training from scratch, another line of research orchestrates pre-trained unimodal experts to leverage strong priors; MMDisCo [13] uses a discriminator for cooperative guidance, while Universe-1 [14] fuses specialized models at the block level. Despite advances in architecture and data scale—with recent works leveraging larger corpora like VGGSound [15] and AudioSet [16]—most models are confined to generating general sound effects or music. The synthesis of synchronized speech alongside video remains a significant and underexplored challenge due to the requirement for fine-grained lip synchronization.
@@ 2.2 Voice Cloning
Voice Cloning aims to mimic a certain speaking style, including timbre, prosody, emotion and more nuanced features. Substantial advancements have been achieved for this task, especially from the scope of TTS (text-to-speech), in a zero-shot way. The conditioning pattern for the voice cloning task are of three general paradigms: (1) providing solely an audio prompt for in-context learning [17, 18, 19, 20, 21, 22], where all features of the given speaker are reserved but deeply entangled and uncontrollable; (2) further leveraging a speaker embedding (or derived from a specific aspect of speaking style, e.g. emotion) to provide overall guidance for synthesis [23, 24, 25, 26], which allows partial controllability within a relatively simple model design; (3) instructing with text prompt depicting desired variations of the cloned voice [27, 28, 29], which offers extremely high flexibility but still underexplored. Although there have been many explorations in pure speech tasks, there have been few studies on cross-domain transition (covering tasks such as music and sound) and multimodal generation in terms of voice cloning.
@@ 3.1 Multimodal Diffusion Transformer
Our model employs a unified conditioning architecture that accommodates diverse input modality configurations—including text, image, and audio—in arbitrary combinations. This design enables the model to handle various generation scenarios, such as T2VA (text-to-video-and-audio), TI2VA (text-and-image-to-video-and-audio), TA2VA (text-and-audio-to-video-and-audio), and TIA2VA (text-image-audio-to-video-and-audio), within a single framework. The proposed approach builds upon the MM-DiT architecture introduced in SD3[30]. During training, each input modality is processed by dedicated encoders, with absent modalities represented by trainable placeholder tokens. After independent encoding, all modality-specific features are projected into a shared embedding space to facilitate cross-modal integration.
@@ 3.2 Flow Matching
Our training framework leverages conditional flow matching [31] as the generative modeling objective [32]. The approach learns a parameterized velocity field $v_{\theta}(t,C,x)$ conditioned on inputs $C$ (such as encoded text or video representations) that characterizes the transformation dynamics at each timestep $t$, where $\theta$ denotes the trainable network weights. We adopt the Optimal Transport(OT) path to construct the flow, defining the ground-truth vector field as $u_{t}(x|x_{1})=x_{1}-(1-t)x_{0}$. This learned field enables the transformation of an initial Gaussian noise sample $x_{0}$ into the target audio latent $x_{1}$ through numerical integration via an ODE solver across the temporal interval $t\in[0,1]$. Here, $p$ indicates the conditional probability trajectory and $q$ represents the empirical data distribution.
@@ 3.3 Audio-Video Latent Representation
Latent Audio Encoder-Decoder. The latent audio codec extends our prior SecoustiCodec framework[33, 34, 35], inheriting its core architecture while introducing key modifications to optimize audio reconstruction quality. The codec employs a Mel-VAE structure comprising three primary components: a mel-spectrogram encoder, a mel-spectrogram decoder, and a discriminator. The audio encoder processes input waveforms sampled at 44.1 kHz and compresses them into continuous latent embeddings at a substantially reduced temporal rate, achieving high compression ratios while preserving perceptual quality. The VAE architecture enables the model to learn continuous and complete distributions in the latent space, significantly enhancing audio representation capabilities and reconstruction fidelity
Latent Video Encoder-Decoder. For visual representation, we adopt the 3D causal temporal encoder from CogVideoX[36] as our video codec backbone. The video encoder processes input frames at a resolution of $H\times W$ and temporal length $T$, compressing them into spatiotemporal latent embeddings with dimensionality $h\times w\times t$, where $h\ll H$, $w\ll W$, and $t\ll T$. The encoder applies progressive spatial-temporal downsampling, achieving a total compression ratio of $(H/h)\times(W/w)\times(T/t)$ to produce compact latent representations. Correspondingly, the video decoder mirrors the encoder architecture with transposed convolutions, reconstructing high-fidelity video frames from the latent embeddings. Similar to the audio codec, the video VAE learns a continuous latent distribution, facilitating smooth interpolation and stable training dynamics in the flow matching framework.
@@ 4.2 Unified Instruction-Guided Input
To enable unified audio-video generation, we design a standardized instruction-phoneme input format that maintains consistent text modality inputs across all tasks, as illustrated in Fig. 1. This format comprises three key components: a video instruction description, an audio instruction description, and a phoneme sequence. The video instruction description encapsulates the complete visual scene information required for generation. The audio instruction description specifies the desired acoustic attributes of the output[37]. For speech synthesis, these attributes include speaker characteristics such as gender, age, emotion, style, and accent. In dialogue scenarios, we provide separate descriptions for each of the two speakers and prepend special tokens [S0] and [S1] to their respective text inputs to differentiate the utterances. Similarly, for music generation, the description defines attributes like the singer’s timbre (e.g., gender and age), music genre, instrumentation, melody, and emotional expression. For sound effect generation, the audio description contains the relevant sonic scene information. In both speech and music tasks, the input text (for speech) or lyrics (for music) is converted into a phoneme sequence using a Grapheme-to-Phoneme (G2P) model [38]. This unified representation allows a single model to seamlessly process text modality inputs for diverse generation tasks. It is important to note that in the T2VA and TI2VA tasks, the timbre of the generated audio is solely controlled by the audio instruction description. However, in the TA2VA and TIA2VA tasks, the timbre is derived from the input reference audio. These multimodal conditional inputs can coexist; for instance, a reference audio input can be accompanied by an audio instruction description such as "this woman speaks happily and surprisingly…".
@@ 4.5 Noise-Based Negative Conditioning
A naive approach to constructing the unconditional acoustic branch is to use a zero vector (i.e., $S=0$) as the negative speaker embedding. However, in our training strategy, a zero vector specifically signals the T2VA task, where the model is trained to generate a random or context-appropriate voice from text instructions alone. Consequently, using a zero vector as the negative condition in CFG merely instructs the model to diverge from the "random voice generation" mode rather than steering it away from low-quality audio. To address this, we propose a Noise-based Negative Speaker Embedding strategy. We construct the negative speaker embedding $S_{neg}$ by feeding noise into the pre-trained speaker encoder, establishing a repulsive force against unstructured acoustic features.
@@ 5.1 High-Fidelity Synthetic Timbre Dataset
Figure 2: The data construction pipeline for the high-fidelity synthetic timbre dataset. The process begins with extracting clean acoustic prompts from raw speech and music via denoising and vocal separation tools. These prompts condition DMP-TTS [29] to synthesize speech from neutral texts. Finally, a WavLM-based[39] speaker verification module filters out samples with low cosine similarity to ensure high speaker identity preservation.
Constructing a dataset for zero-shot voice cloning requires decoupling speaker timbre from other acoustic factors (e.g., emotion, background noise) while maintaining high identity similarity. As illustrated in Figure 2, we implemented a rigorous synthetic data construction pipeline using DMP-TTS [29], a zero-shot text-to-speech model.
First, to obtain high-quality acoustic prompts, we sampled raw speech and music data from our internal repository. We utilized proprietary source separation tools to denoise raw speech and extract dry vocals from musical tracks, ensuring the reference timbre was free from environmental artifacts. These processed segments served as the acoustic conditions. Second, to eliminate emotional leakage and ensure stability, we synthesized speech based on a pre-defined set of emotionally neutral texts, explicitly instructing the TTS model to generate neutral-style speech. Finally, to guarantee the preservation of speaker identity, we employed a speaker verification filter based on WavLM-large [39]. We calculated the cosine similarity between the speaker embeddings of the synthesized speech and the ground truth prompts, discarding any samples that failed to meet a strict similarity threshold. This process yielded a high-quality dataset specifically optimized for learning timbre representations.
@@ 6 Experiments
AudioLDM+TemoTkn[40, 41] Cascaded 0.05 0.12 0.07 2.05 2.53 0.229 1.15 2.68 0.137
OpenSora+FoleyGen[42, 43] 0.42 0.44 0.56 3.69 3.08 0.212 0.92 2.76 0.159
OpenSora+See&Hear[42, 44] 0.42 0.44 0.56 2.26 2.97 0.206 0.337 0.86 2.85 0.164
JavisDiT[8] Unified 0.18 0.36 0.22 1.95 5.17 0.228 0.256 0.92 3.94 0.231
SVG[45] 0.40 0.41 0.25 1.55 3.62 0.080 0.72 4.07 0.206
Universe-1[14] 0.20 0.47 0.25 1.55 1.25 0.160 0.180 0.98 3.92 0.198
Ovi[9] 0.58 0.48 0.46 1.50 1.19 0.224 0.035 0.82 4.28 0.214
@@ 6.1 Implementation Details
MM-Sonate is built with 20 billion parameters and utilizes a flow matching feedforward dimension of 4096. Its backbone consists of 32 joint diffusion transformer layers equipped with RoPE positional encoding [46]. For text conditioning, we adopt a phoneme encoder with a feedforward dimension of 512, alongside Qwen2.5-7B [47] serving as the caption encoder. The video VAE accommodates inputs of varying resolutions and frame rates, applying 16× spatial compression (height and width) to produce embeddings at a temporal rate of 3 Hz. Furthermore, audio processing is handled by a mel encoder that downsamples 44.1 kHz waveforms by a factor of 1024, yielding embeddings at 43 Hz.
@@ 6.2 Compared Method
To comprehensively evaluate MM-Sonate, we benchmark against SOTA models across three distinct tasks. Notably, since existing T2VA models do not support timbre cloning capabilities, we compare MM-Sonate with TTS models. It is important to note that all TTS and TTM comparisons for MM-Sonate simultaneously generate accompanying video content. For audio-video joint generation, we compare against both cascade systems, including AudioLDM+TemoTkn [40, 41], OpenSora+FoleyGen [42, 43] and OpenSora+See&Hear [42, 44], and unified models, such as JavisDiT [8], SVG [45], Universe-1 [14], and the current open-source state-of-the-art model Ovi [9]. For TTS task, we compare against MaskGCT [48], E2-TTS [49], F5-TTS [18], ZipVoice [50], M3-TTS [22], and CosyVoice2 [28]. For TTM task, we compare against DiffRhythm+[51] and ACE-Step[52]. All evaluations are conducted on VerseBench, the benchmark dataset released with Universe-1 [14].

### v2

@@ Abstract
Figure 1: Audio–Video Joint Generation with Multimodal Control: synthesizing coherent, temporally aligned dialogue conditioned on the first-frame and distinct reference audios. Unlike existing frameworks, our unified architecture supports flexible combinations of multi-modal inputs for precise control. MM-Sonate outperforms baselines (Ovi, CosyVoice2) on 9/11 metrics.
@@ 1 Introduction
The ultimate goal of generative world simulation is to replicate the multisensory nature of physical reality, necessitating the synchronous synthesis of visual and acoustic signals. While recent text-to-video diffusion models have achieved remarkable visual fidelity, the integration of high-quality, synchronized audio remains a significant challenge. Early approaches typically treated audio generation as a post-processing step, employing cascaded systems where audio models are conditioned on pre-generated video frames Cheng et al. (2024); Wang et al. (2025c); Luo et al. (2023); Wang et al. (2025d); Guo et al. (2025b; a). However, this decoupled paradigm fundamentally struggles with fine-grained temporal alignment, frequently resulting in significant asynchrony between video and audio streams.
To address the alignment problem, research has shifted toward unified architectures that generate audio and video within a single model. Pioneering works such as MM-Diffusion Ruan et al. (2023) and JavisDiT Liu et al. (2025) explored joint modeling, while state-of-the-art models like Ovi Low et al. (2025) achieve impressive synchronization through symmetric dual-backbone designs. Despite these advances, a critical limitation persists: existing unified models lack fine-grained controllability over the acoustic modality, particularly in speech generation. For instance, while Ovi can generate semantically relevant sounds, it cannot control the specific identity (timbre) of a speaker or clone a voice from a reference clip. Consequently, current unified video generators cannot yet function as high-fidelity, personalized speech synthesis, limiting their application in scenarios requiring consistent character identity or dubbing.
@@
22footnotetext: “Sonate” is derived from the musical term “Sonata”, symbolizing the model’s comprehensive capabilities in audio generation.
@@ 2.1 Audio-Video Joint Generation
Audio-Video Joint Generation aims to simultaneously synthesize visual frames and acoustic signals, typically conditioned on text descriptions, ensuring both semantic consistency and temporal synchronization across modalities within a unified generation process. Pioneering efforts like MM-Diffusion Ruan et al. (2023) introduced this task using coupled U-Net backbones but were often constrained by small-scale datasets (e.g., roughly 10 hours), which limited their generalization capabilities. Following the architectural shift in unimodal generation, recent works have predominantly adopted Diffusion Transformers (DiTs). These approaches generally fall into two architectural paradigms: unified or coupled streams. For instance, UniForm Zhao et al. (2025) employs a unified single-tower architecture to process concatenated audio-video tokens, whereas AV-DiT Wang et al. (2024a) adapts a pre-trained image DiT with lightweight adapters to handle multimodal signals. A central challenge in this field remains precise spatio-temporal synchronization. To address this, JavisDiT Liu et al. (2025) introduces a hierarchical prior, while Ovi Low et al. (2025) enforces tight coupling through a symmetric twin-backbone design. Similarly, SyncFlow Liu et al. (2024) utilizes a dual-DiT architecture with Rectified Flow Matching to enhance alignment stability. Apart from training from scratch, another line of research orchestrates pre-trained unimodal experts to leverage strong priors; MMDisCo Hayakawa et al. (2024) uses a discriminator for cooperative guidance, while Universe-1 Wang et al. (2025a) fuses specialized models at the block level. Despite advances in architecture and data scale—with recent works leveraging larger corpora like VGGSound Chen et al. (2020) and AudioSet Gemmeke et al. (2017)—most models are confined to generating general sound effects or music. The synthesis of synchronized speech alongside video remains a significant and underexplored challenge due to the requirement for fine-grained lip synchronization.
@@ 2.2 Voice Cloning
Voice Cloning aims to mimic a certain speaking style, including timbre, prosody, emotion and more nuanced features. Substantial advancements have been achieved for this task, especially from the scope of TTS (text-to-speech), in a zero-shot way. The conditioning pattern for the voice cloning task are of three general paradigms: (1) providing solely an audio prompt for in-context learning Chen et al. (2025b; 2024); Qiang et al. (2024a; c); Wang et al. (2025b; e), where all features of the given speaker are reserved but deeply entangled and uncontrollable; (2) further leveraging a speaker embedding (or derived from a specific aspect of speaking style, e.g. emotion) to provide overall guidance for synthesis Du et al. (2024a); Cui et al. (2025); Zhou et al. (2025); Yu et al. (2025), which allows partial controllability within a relatively simple model design; (3) instructing with text prompt depicting desired variations of the cloned voice Ji et al. (2024); Du et al. (2024b); Yin et al. (2025), which offers extremely high flexibility but still underexplored. Although there have been many explorations in pure speech tasks, there have been few studies on cross-domain transition (covering tasks such as music and sound) and multimodal generation in terms of voice cloning.
@@ 3.1 Multimodal Diffusion Transformer
Our model employs a unified conditioning architecture that accommodates diverse input modality configurations—including text, image, and audio—in arbitrary combinations. This design enables the model to handle various generation scenarios, such as T2VA (text-to-video-and-audio), TI2VA (text-and-image-to-video-and-audio), TA2VA (text-and-audio-to-video-and-audio), and TIA2VA (text-image-audio-to-video-and-audio), within a single framework. The proposed approach builds upon the MM-DiT architecture introduced in SD3Esser et al. (2024). During training, each input modality is processed by dedicated encoders, with absent modalities represented by trainable placeholder tokens. After independent encoding, all modality-specific features are projected into a shared embedding space to facilitate cross-modal integration.
@@ 3.2 Flow Matching
Our training framework leverages conditional flow matching Lipman et al. (2022) as the generative modeling objective Tong et al. (2023). The approach learns a parameterized velocity field $v_{\theta}(t,C,x)$ conditioned on inputs $C$ (such as encoded text or video representations) that characterizes the transformation dynamics at each timestep $t$, where $\theta$ denotes the trainable network weights. We adopt the Optimal Transport(OT) path to construct the flow, defining the ground-truth vector field as $u_{t}(x|x_{1})=x_{1}-(1-t)x_{0}$. This learned field enables the transformation of an initial Gaussian noise sample $x_{0}$ into the target audio latent $x_{1}$ through numerical integration via an ODE solver across the temporal interval $t\in[0,1]$. Here, $p$ indicates the conditional probability trajectory and $q$ represents the empirical data distribution.
@@ 3.3 Audio-Video Latent Representation
Latent Audio Encoder-Decoder. The latent audio codec extends our prior SecoustiCodec frameworkQiang et al. (2025b; a; 2024b), inheriting its core architecture while introducing key modifications to optimize audio reconstruction quality. The codec employs a Mel-VAE structure comprising three primary components: a mel-spectrogram encoder, a mel-spectrogram decoder, and a discriminator. The audio encoder processes input waveforms sampled at 44.1 kHz and compresses them into continuous latent embeddings at a substantially reduced temporal rate, achieving high compression ratios while preserving perceptual quality. The VAE architecture enables the model to learn continuous and complete distributions in the latent space, significantly enhancing audio representation capabilities and reconstruction fidelity
Latent Video Encoder-Decoder. For visual representation, we adopt the 3D causal temporal encoder from CogVideoXYang et al. (2024) as our video codec backbone. The video encoder processes input frames at a resolution of $H\times W$ and temporal length $T$, compressing them into spatiotemporal latent embeddings with dimensionality $h\times w\times t$, where $h\ll H$, $w\ll W$, and $t\ll T$. The encoder applies progressive spatial-temporal downsampling, achieving a total compression ratio of $(H/h)\times(W/w)\times(T/t)$ to produce compact latent representations. Correspondingly, the video decoder mirrors the encoder architecture with transposed convolutions, reconstructing high-fidelity video frames from the latent embeddings. Similar to the audio codec, the video VAE learns a continuous latent distribution, facilitating smooth interpolation and stable training dynamics in the flow matching framework.
@@ 4.2 Unified Instruction-Guided Input
To enable unified audio-video generation, we design a standardized instruction-phoneme input format that maintains consistent text modality inputs across all tasks, as illustrated in Fig. 2. This format comprises three key components: a video instruction description, an audio instruction description, and a phoneme sequence. The video instruction description encapsulates the complete visual scene information required for generation. The audio instruction description specifies the desired acoustic attributes of the outputQiang et al. (2025c). For speech synthesis, these attributes include speaker characteristics such as gender, age, emotion, style, and accent. In dialogue scenarios, we provide separate descriptions for each of the two speakers and prepend special tokens [S0] and [S1] to their respective text inputs to differentiate the utterances. Similarly, for music generation, the description defines attributes like the singer’s timbre (e.g., gender and age), music genre, instrumentation, melody, and emotional expression. For sound effect generation, the audio description contains the relevant sonic scene information. In both speech and music tasks, the input text (for speech) or lyrics (for music) is converted into a phoneme sequence using a Grapheme-to-Phoneme (G2P) model Qiang et al. (2022). This unified representation allows a single model to seamlessly process text modality inputs for diverse generation tasks. It is important to note that in the T2VA and TI2VA tasks, the timbre of the generated audio is solely controlled by the audio instruction description. However, in the TA2VA and TIA2VA tasks, the timbre is derived from the input reference audio. These multimodal conditional inputs can coexist; for instance, a reference audio input can be accompanied by an audio instruction description such as ”this woman speaks happily and surprisingly…”.
@@ 4.5 Noise-Based Negative Conditioning
A naive approach to constructing the unconditional acoustic branch is to use a zero vector (i.e., $S=0$) as the negative speaker embedding. However, in our training strategy, a zero vector specifically signals the T2VA task, where the model is trained to generate a random or context-appropriate voice from text instructions alone. Consequently, using a zero vector as the negative condition in CFG merely instructs the model to diverge from the ”random voice generation” mode rather than steering it away from low-quality audio. To address this, we propose a Noise-based Negative Speaker Embedding strategy. We construct the negative speaker embedding $S_{neg}$ by feeding noise into the pre-trained speaker encoder, establishing a repulsive force against unstructured acoustic features.
@@ 5.1 High-Fidelity Synthetic Timbre Dataset
Figure 3: The data construction pipeline for the high-fidelity synthetic timbre dataset. The process begins with extracting clean acoustic prompts from raw speech and music via denoising and vocal separation tools. These prompts condition DMP-TTS Yin et al. (2025) to synthesize speech from neutral texts. Finally, a WavLM-basedChen et al. (2022) speaker verification module filters out samples with low cosine similarity to ensure high speaker identity preservation.
Constructing a dataset for zero-shot voice cloning requires decoupling speaker timbre from other acoustic factors (e.g., emotion, background noise) while maintaining high identity similarity. As illustrated in Figure 3, we implemented a rigorous synthetic data construction pipeline using DMP-TTS Yin et al. (2025), a zero-shot text-to-speech model.
First, to obtain high-quality acoustic prompts, we sampled raw speech and music data from our internal repository. We utilized proprietary source separation tools to denoise raw speech and extract dry vocals from musical tracks, ensuring the reference timbre was free from environmental artifacts. These processed segments served as the acoustic conditions. Second, to eliminate emotional leakage and ensure stability, we synthesized speech based on a pre-defined set of emotionally neutral texts, explicitly instructing the TTS model to generate neutral-style speech. Finally, to guarantee the preservation of speaker identity, we employed a speaker verification filter based on WavLM-large Chen et al. (2022). We calculated the cosine similarity between the speaker embeddings of the synthesized speech and the ground truth prompts, discarding any samples that failed to meet a strict similarity threshold. This process yielded a high-quality dataset specifically optimized for learning timbre representations.
@@ 6 Experiments
AudioLDM+TemoTknLiu et al. (2023); Yariv et al. (2024) Cascaded 0.05 0.12 0.07 2.05 2.53 0.229 1.15 2.68 0.137
OpenSora+FoleyGenOpenAI (2024); Mei et al. (2024) 0.42 0.44 0.56 3.69 3.08 0.212 0.92 2.76 0.159
OpenSora+See&HearOpenAI (2024); Xing et al. (2024) 0.42 0.44 0.56 2.26 2.97 0.206 0.337 0.86 2.85 0.164
JavisDiTLiu et al. (2025) Unified 0.18 0.36 0.22 1.95 5.17 0.228 0.256 0.92 3.94 0.231
SVGIshii et al. (2024) 0.40 0.41 0.25 1.55 3.62 0.080 0.72 4.07 0.206
Universe-1Wang et al. (2025a) 0.20 0.47 0.25 1.55 1.25 0.160 0.180 0.98 3.92 0.198
OviLow et al. (2025) 0.58 0.48 0.46 1.50 1.19 0.224 0.035 0.82 4.28 0.214
@@ 6.1 Implementation Details
MM-Sonate is built with 20 billion parameters and utilizes a flow matching feedforward dimension of 4096. Its backbone consists of 32 joint diffusion transformer layers equipped with RoPE positional encoding Su et al. (2024). For text conditioning, we adopt a phoneme encoder with a feedforward dimension of 512, alongside Qwen2.5-7B Yang et al. (2025) serving as the caption encoder. The video VAE accommodates inputs of varying resolutions and frame rates, applying 16× spatial compression (height and width) to produce embeddings at a temporal rate of 3 Hz. Furthermore, audio processing is handled by a mel encoder that downsamples 44.1 kHz waveforms by a factor of 1024, yielding embeddings at 43 Hz.
@@ 6.2 Compared Method
To comprehensively evaluate MM-Sonate, we benchmark against SOTA models across three distinct tasks. Notably, since existing T2VA models do not support timbre cloning capabilities, we compare MM-Sonate with TTS models. It is important to note that all TTS and TTM comparisons for MM-Sonate simultaneously generate accompanying video content. For audio-video joint generation, we compare against both cascade systems, including AudioLDM+TemoTkn Liu et al. (2023); Yariv et al. (2024), OpenSora+FoleyGen OpenAI (2024); Mei et al. (2024) and OpenSora+See&Hear OpenAI (2024); Xing et al. (2024), and unified models, such as JavisDiT Liu et al. (2025), SVG Ishii et al. (2024), Universe-1 Wang et al. (2025a), and the current open-source state-of-the-art model Ovi Low et al. (2025). For TTS task, we compare against MaskGCT Wang et al. (2024b), E2-TTS Eskimez et al. (2024), F5-TTS Chen et al. (2024), ZipVoice Zhu et al. (2025), M3-TTS Wang et al. (2025e), and CosyVoice2 Du et al. (2024b). For TTM task, we compare against DiffRhythm+Chen et al. (2025a) and ACE-StepGong et al. (2025). All evaluations are conducted on VerseBench, the benchmark dataset released with Universe-1 Wang et al. (2025a).